Home‎ > ‎Crew 162‎ > ‎

0130-DailySummaryReport

Crew 162 Daily Summary Report for Sol 07 01/30/2016

Summary Title: Transmissions to Earth

Mission Status: Half-way point reached.

Sol Activity Summary: Crew 162 worked on chores, research and education
& outreach video messages for transmission back to Earth. We continued
to troubleshoot the 3d printer and were able to create a Tintin style
moon rocket that our HSO added stars to in the 3d modeling program. We
also had a lesson in navigational geography this morning while
troubleshooting our GPS unit and comparing coordinates against the paper
maps we have a the hab. For EVA planning here at MDRS, we utilize the
NAD27 coordinate system which differs from the coordinates used by
world-wide Earth GPS units. It takes some getting used to but it's a
very efficient system comprised of northing and easting measurements.

Look Ahead Plan: Severe winter weather beginning tonight, possible snow
on Sunday Sol 08. We will have our second media visit on Tuesday, Sol
10, by a professional photographer from London.

Anomalies in work: No current anomalies to report. We are still locked
out of the observatory but astronomy mission support is working to get a
replacement padlock and key.

Weather: Dark clouds rolled in prior to sunset, as well as high winds.
Wet rain began after sundown.

Crew Physical Status: Crew is sleeping more but reporting feeling tired
towards the end of each day. Otherwise crew is in good physical health.

EVA: EVA #3 completed by HSO and Crew Engineer. Report to be submitted
tonight.

Reports to be filed:
Commanders Report
Engineering Report
Journalist Report
GreenHab Report
EVA #3 Report
Daily Photos

Support Requested: No additional support requested from Mission Control
at this time.